Consider a voltage amplifier having a frequency response of the low-

pass STC type with a dc gain of 60 dB and a 3-dB frequency of 1000 

Hz. Find the gain in dB at f = 10 Hz


1
Expert's answer
2021-02-25T10:53:18-0500

First, we will use the

"|T(j\omega)| = \large\frac{|K|}{\sqrt{1+(\large\frac{\omega}{\omega_o})^2}}"

low pass Magnitude response equation.

"A_{VdB} = 20*log_{10}(A_V)"

"60 = 20*log_{10}(A_{V})"

"3 = log_{10}(A_V) \to A_V = 10^3 = 1000"

Next, knowing that "\large\frac{f}{f_o} = \frac{\omega}{\omega_o}", and that "f_o" is given we can plug that in formula:

"A_V(f) = \frac{1000}{\sqrt{1+(\frac{f}{1000})^2}} =" "995.95"